OpenStudy (okdutchman7):

Please help. WILL MEDAL! @Mehek14 @nincompoop @sleepyjess @KendrickLamar2014 @satellite73 @zepdrix @Loser66 @Elsa213 @xapproachesinfinity @Owlcoffee

OpenStudy (okdutchman7):

\[2 \ln 1 + 2 \ln 5 = \ln (x+3)\]

OpenStudy (loser66):

I saw you post it more than once. Why?

OpenStudy (okdutchman7):

\[\ln 1 + \ln 25 = \ln (x+3)\]

OpenStudy (okdutchman7):

I don't think I posted this one before.

OpenStudy (okdutchman7):

Oh well that explain it then.